IOP6 is slated to begin this evening when the radar crews get their radars running, but no later that 10 PM PST. Zonal flow today brings a nice surge of vapor on west winds increasing to 25-30 kt overnight. Precip water amounts are over an inch during the precipitation period, which appears to start after dark and continue until midday tomorrow according to the NAM model - the GFS model holds on to precip until sundown. Snow level starts out around 9000 ft, lowers to around 7500 ft by the end of the event. Total model precipitation amounts .5 to 1.0 inches.
Supplementary soundings from OAK are slated to start at 03Z Wednesday and continue to 21 Z Wednesday. The Slough House sounding site will launch 3 hourlies starting at 00Z Wednesday to 00Z Thursday, but may be continued one more cycle at 03Z if students can do it and the precip holds out.
Next event starts Friday afternoon. Sharp, fast-moving trough moves in, strong warm advection ahead of it, 850mb winds southerly at 35+ kt bring precipitable water amounts around 1 inch. Precipitation amounts not large (.5 inch) before warm advection gives way to cold advection after dark Friday. Strong cold advection and frontogenesis in predawn hours appear to cause greatest precipitation rates, 1 inch in 6 hrs ending sunrise. Total precip 1.5 to 2 inches, event ends late Saturday night.
The next event following the Saturday storm looks to start the following Tuesday morning (Jan. 17). There will be a change of crews for the SMART-R following the Friday/Saturday event and new crews are alerted for possbile departure on Monday for operations beginning Tuesday morning.
Status of HMT Instrument Problems:
The New Years' storms caused problems for various HMT instruments and we are working hard to recover from that hit. Some, such as the surface met stations at Chico and Slough House, have recently been fixed and are back on line. Many of the problems are associated with the communications links; data are being recorded but not displayed in realtime. Other instruments have comm link problems that pre-date the New Years storms. As of this morning, the following HMT sites do not have realtime display of data on the project Web page, but all are being worked on, as shown.
BBY - 915, 449 profilers, sfc met (no realtime display since 1JAN06);
a new phone line is supposed to be installed there early this week.
SHS - 915 profiler (down since 1JAN06);
Jesse Leach will install the repaired transmitter there on Friday.
CFC - S-band profiler, sfc met, disdrometer, hotplate (no realtime display since 6JAN06);
Clark King will be there Wednesday to assess the problem.
PPB - 915 profiler, sfc met (has never been available for realtime display this winter);
Scott Abbott has recovered from surgery and will install a new satellite-comm system there on Wednesday.
FHL - sfc met (has never been available for realtime display this winter);
the small local phone company must install a phone line - we keep pushing them to do it.
As soon as data from any of these sites starts flowing into Boulder, Dave White will get them into the data system and available for realtime display through the HMT-06 web page.